chloew3:

Mickey deposited $782.11 in a savings account that earns 4.3% simple interest. What is Mickey's account balance after 16 years?

Narad:

The interest is =\[782.11*4.3*16/100=538.09 $\] The account balance after 16 years is \[=782.11+538.09 = 1320.20$\]

fwval:

e